### v1

# THOMAS et al

Abstract
[Summary]We make use of Kronecker structure for scaling Gaussian Process models to large-scale, heterogeneous, clinical data sets. Repeated measures, commonly performed in clinical research, facilitate computational acceleration for nonlinear Bayesian nonparametric models and enable exact sampling for non-conjugate inference, when combinations of continuous and discrete endpoints are observed. Model inference is performed in Stan, and comparisons are made with brms on simulated data and two real clinical data sets, following a radiological image quality theme. Scalable Gaussian Process models compare favourably with parametric models on real data sets with 17,460 observations. Different GP model specifications are explored, with components analogous to random effects, and their theoretical properties are described.
[…]
keywords:
Gaussian Processes, Scalable Inference, Repeated Measures, Kronecker Products, Heterogeneous Multi-output Regression, Stan, Radiological Image Quality
articletype: Article Type00footnotetext: Abbreviations: GP, Gaussian Process; HMC, Hamiltonian Monte Carlo; RCT, Randomised Control Trial; ARD, Automatic Relevance Determination; NUTS, No U-Turn Sampler; CTA, Computed Tomography Angiography; keV, kiloelectronVolts; ROI, Region Of Interest; SVM, Support Vector Machine; ESS, Effective Sample Size; HU, Hounsfield Unit; ICM, Intrinsic Coregionalization Model
1 Introduction
Clinical research is often performed with structured data built into the study design, sometimes by repeated measurements of individuals at different time points or locations, or using different measurements methods simultaneously on the same individuals. For example, longitudinal studies follow the same cohort repeatedly at different points using the same measurement process, while many Randomised Control Trials (RCTs) will simultaneously measure primary and secondary endpoints reflecting different aspects of a clinical process.
In the article, we observe that either of these types of structured measurement can be used to enable computational tractability of a class of complex statistical models that would otherwise not scale to real clinical data sets of thousands of measurements. Specifically, appropriate repeated measurements enable the pursuit of exact inference for Gaussian Process (GP)williams2006gaussian models by representing their covariance matrices as Kronecker productssaatcci2012scalable. GPs are Bayesian nonparametric models, that are capable of capturing nonlinear covariate dependence, or multi-output correlationsalvarez2012kernels that would ordinarily be neglected by commonly-used parametric statistical models.
Here we present and run new GP models using repeated measurements in the covariate structure and multiple heterogeneous (mixed continuous and discrete) outputs: exact inference in Stanstancore is scaled to tens of thousands of measurements on the processors of a domestic-issue laptop. The models represent the joint covariance between all of the data as Kronecker-structured, and perform Hamiltonian Monte Carlo (HMC) sampling for hyperparameters, missing output values, and on the latent GP space for non-conjugate inference in the presence of heterogeneous outputs. The clinical data sets follow the theme of radiological image quality, in which the covariates consist of patient characteristics, body locations and time, while the outputs represent measurements of image quality, either continuously "objectively" on the Hounsfield scalehounsfieldunits of radiodensity, or discretely "subjectively" from expert evaluations.
{boxtext}
Main Article Contributions
  • The observation that widely-used repeated measurements in clinical research, either in the form of making the same measurement at different points, or making different measurements simultaneously, facilitates the use of Kronecker structure in the covariance matrix of Gaussian Process models and thereby the scaling of exact sampling for Bayesian nonparametric models to large clinical data sets with modest computational resources.
  • An implementation of the relevant Gaussian Process models in Stanstancore, rstan (with wrapper in R), allowing for running the inference on multiple real world data sets, and benchmarking various model specificatons such as covariance function choices.
  • An empirical investigation using real-world clinical data sets into the predictive abilities of the Gaussian Process models compared to standard parametric regression models run in brmsbrms, demonstrating the utility of more complex models.
2 Methods
In this section, we describe the methodological concepts relevant to the models implemented in this article.
2.1 Repeated Measurements
The value of repeated measurements is well-established and widely-understood in clinical research. The use of random effects is common when data is drawn from a structured population for which a hierarchical model is appropriate, for example when measuring different individuals repeatedly, or in a multi-centre study. Further structure emerges if data is collected in a systematic way, for example at consistent follow-up times for the entire population, or at multiple, consistent anatomical locations in the body for scans or biopsies. Data can also exhibit repetitions at the outcome level, when different outcomes of interest are measured at the same locations, individuals, and time points: this is common in RCTs with a combination of primary and secondary endpoints.
Within a regression framework, the multiple outcomes can be represented as a matrix Y, and the covariates corresponding to treatments, locations, times, patient characteristics, or anything else that might influence the outcomes, can be represented as a covariate matrix X. Different clinical study designs will imposed different structure on the covariate matrix X. If we consider a design in which $N_{1}$ individuals are measured at $N_{2}$ time points, at $N_{3}$ anatomical locations, with the time points and anatomical locations being identical between individuals, then we can divide columnwise the long-format covariate matrix X of height $N_{1}N_{2}N_{3}$ into matrices $X_{1}$, $X_{2}$, and $X_{3}$, with $X_{1}$ containing information $x_{1}$ about the $N_{1}$ unique individuals, repeated $N_{2}N_{3}$ times, $X_{2}$ containing information $x_{2}$ about the $N_{2}$ unique times of measurement, repeated $N_{1}N_{3}$ times, and $X_{3}$ containing information $x_{3}$ about the $N_{3}$ unique anatomical locations, repeated $N_{1}N_{2}$ times. There are in addition $N_{4}$ distinct outcomes measured for every value of $X$, resulting in an outcome matrix Y of size $N_{1}N_{2}N_{3}$ by $N_{4}$.
2.2 Gaussian Processes
Gaussian Processes (GPs) are nonlinear, Bayesian models designed for flexible, probabilistic supervised learning williams2006gaussian. They model a dependent variable $y$ conditional on independent variables $X$, via a mean function $m(x)$ and covariance111also known as kernel function function $k(x,x^{\prime},\theta)$, defining a latent variable $f$ that is joint-normally distributed over all the observed data points. The covariance functions are described by hyperparameters $\theta$ that can be learned from data, while the mean function will be taken to be zero here with no loss of generality. The latent function $f$ can be passed through a Gaussian likelihood with a noise variance $\sigma^{2}$ to model a continuous output $y$:
$f\sim$ $\mathcal{GP}(0,k(x,x^{\prime},\theta))$ (1)
$y|x=$ $f(x)+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2})$ (2)
Specific choices of covariance functions $k(x,x^{\prime},\theta)$ have corresponding implicit parametric basis functions. One advantage of the "function-space" formulation is the ability to use a finite representation of a function with a potentially infinite-dimensional parametric representation: this is the sense in which the models are considered "nonparametric", in that they avoid specifying a parametric model for the latent mean function.
2.3 Scalability and Kronecker-Structure Covariance Matrices
Asserting a GP with a covariance function $k(x,x^{\prime},\theta)$ over a data set with $N$ covariate observations defines a $N\times N$ covariance222also known as kernel matrix matrix $\mathbf{K}$, where the matrix element $\mathbf{K}[i,j]$ is equal to the covariance function $k(x^{i},{x^{j}}^{\prime},\theta)$ evaluated at the $ith$ and $jth$ data point. One challenge of working with GPs is the need to perform a decomposition of the $N\times N$ covariance matrix, which is the dominant computational demand when evaluating the marginal likelihood for sampling or computing the predictive distributions. This results in $\mathcal{O}(N^{3})$ cubic scaling in computational costs with the number of data points $N$, ruling out exact inference for generic covariance matrices for larger data sets . Various methods exist to enable approximate inference for larger data setsliu2020gaussian, thomas2017scalable, while exact inference is possible for larger data sets when there is structure in the covariance matrix that can be exploited, such as Kronecker or Toeplitzsaatcci2012scalable. In instances where data can be represented as lying in a grid structure, with separable covariance structure between each dimension of the grid, the full covariance matrix can be represented as a Kronecker product between the dimensions. A Kronecker product is an operation performed on two matrices generating a third matrix composed of each individual element of the first matrix separately multiplied with the entire second matrix, combined in a blockwise fashion, i.e. for two matrices $A$ and $B$:
$\mathbf{A}\otimes\mathbf{B}=\begin{bmatrix}a_{11}\mathbf{B}&\cdots&a_{1n}\mathbf{B}\\ \vdots&\ddots&\vdots\\ a_{n1}\mathbf{B}&\cdots&a_{nn}\mathbf{B}\end{bmatrix}$ (4)
For a data set consisting of covariates measured over a two dimensional grid, for example the pixels of an image, the covariance matrix over the entire data set ($\mathbf{K}_{12}=K([x_{1},x_{2}],[x^{\prime}_{1},x^{\prime}_{2}])$) can be represented as the Kronecker product between two covariance matrices representing each grid dimension independently ($\mathbf{K}_{1}=K(x_{1}.x^{\prime}_{1})$, $\mathbf{K}_{2}=K(x_{2}.x^{\prime}_{2}),$), with the constraint that the covariance functions used are separable, i.e. for a grid composed of $n_{1}$ vertical grid points $x_{1}$, and $n_{2}$ horizontal grid points $x_{2}$, then:
$\mathbf{K}_{12}=\mathbf{K}_{1}\otimes\mathbf{K}_{2}.$ (5)
We can apply this reasoning to clinical scenario described in 2.1: for a longitudinal study consisting of $N_{1}$ individuals measured at $N_{2}$ time points at $N_{3}$ anatomical locations, the covariate data X can be represented as lying on a three dimensional grid, where one of the grid dimensions $x_{1}$ is the variation between individuals, the second $x_{2}$ as the variation in time, and the third $x_{3}$ variation in anatomical location. Consequently, defining separate covariance matrices representing the variation across individuals (the $N_{1}\times N_{1}$ matrix $\mathbf{K}_{1}$), variation over time (the $N_{2}\times N_{1}$ matrix $\mathbf{K}_{2}$), and anatomical variation (the $N_{3}\times N_{3}$ matrix $\mathbf{K}_{3}$) we can represent the covariance between the covariate measurements $X$ as:
$\mathbf{K}_{X}=\mathbf{K}_{1}\otimes\mathbf{K}_{2}\otimes\mathbf{K}_{3}.$ (6)
Conveniently, when a decomposition of the full matrix is required, the components of the Kronecker product can be decomposed separately, such that the $\mathcal{O}(N^{3})=\mathcal{O}(N_{1}^{3}N_{2}^{3}N_{3}^{3})$ computational demands become $\mathcal{O}(N_{1}^{3}+N_{2}^{3}+N_{3}^{3})$. For datasets with appropriate grid structure or repeated measurements, this enables exact inference for tens of thousands of data points on a personal computer. This has been used previously for image data, where the pixels lie on a regularly spaced grid, or when there are multiple endpoints evaluated at the same covariate locations. In this article, we note that the widespread use of repeated measurements in clinical research, can be represented within the Kronecker structure described above.
2.4 Heterogeneous Multi-Task Gaussian Processes
The GP framework extends naturally to multi-dimensional responses $\mathbf{Y}$, analogous to the case of multivariate regressionalvarez2012kernels, bonilla2007multi. The different response dimensions (also known as "tasks" or "outputs") are appended into a single vector, and a between-task covariance matrix $\mathbf{K}_{out}$ is used to model correlations between the tasks. When the different tasks are evaluated at the same values of the covariates, generating a covariance matrix between covariates $\mathbf{K}_{X}$ then Kronecker structure again emerges in the full covariance matrix $\mathbf{K}_{f}=\mathbf{K}_{out}\otimes\mathbf{K}_{X}$, and computational accelerations become possible.
Different tasks often correspond to different methods of evaluating some outcome, and often exhibit heterogeneity of distributions, e.g. each task may be variously continuous valued, binary or follow some other distribution. In this case, each output will require different likelihood functions to map from the latent function $f$ to observation $y$. For likelihoods other than Gaussian, the latent variable cannot be integrated out analytically and inference must be performed for the latent variable, through sampling, variational inference, or another approximation.
All of the data sets used in this article use one Kronecker component for the multi-output correlation, and three Kronecker grid components for the covariates, which are grouped into grid dimensions $x_{1}$, $x_{2}$ and $x_{3}$. The full covariance structure is therefore:
$\mathbf{K}_{f}=\mathbf{K}_{out}\otimes\mathbf{K}_{1}\otimes\mathbf{K}_{2}\otimes\mathbf{K}_{3}$ (7)
2.5 Covariance Matrix Design and Random Effects
Imposing Kronecker structure on the full covariance matrix puts some constraints on the types of covariance functions that can be used. Principally, the covariance function must be separable between the different grid components $x_{1}$ and $x_{2}$, meaning the overall covariance function can be represented as the product of covariance functions defined over of the grid components, i.e. $k([x_{1},x_{2}],[x^{\prime}_{1},x^{\prime}_{2}])=k_{1}(x_{1},x^{\prime}_{1})k_{2}(x_{2},x^{\prime}_{2})$. Many commonly used multi-dimensional covariance functions exhibit this property, but some designs that might be desirable for interpretation do not exhibit separability. Two are discussed below: additive covariance functions and random effects covariance functions.
An additive covariance function represents a common covariance function over different dimensions as a sum of separate covariance functionsduvenaud2014automatic, i.e. $k([x_{1},x_{2}],[x^{\prime}_{1},x^{\prime}_{2}])=k_{1}(x_{1},x^{\prime}_{1})+k_{2}(x_{2},x^{\prime}_{2})$. Additive covariance functions might be desired if we are interested in isolating the particular contribution to the output variation from one particular dimension. This is especially useful in medical research when one intervention or treatment is considered to be of central clinical relevance. Within the Kronecker framework, it is possible to assert an additive covariance function within each component of the Kronecker product, but the resulting covariance function over the entire data set is a product of the sum within the component with the covariance functions over the rest of the dimensions, making it different to interpret the sum components separately, i.e.:
$(\mathbf{K_{1}}+\mathbf{K_{2}})\otimes\mathbf{K}_{3}=(\mathbf{K}_{1}\otimes\mathbf{K_{3}})+(\mathbf{K}_{2}\otimes\mathbf{K_{3}})$ (8)
Random effects are often desirable in the presence of repeated measurements, where we are not necessarily interested in the variation between individuals but we would still like to include it in the model gelman2007data. This can be performed in an elegant way within GP regression models by including a structured diagonal noise component representing the random variation across individuals. In the Kronecker context, this can achieved by adding spherical noise to the Kronecker product component that represents the variation between individuals, representing the "noise" sampled when moving between individuals. As this is an additive covariance function where one of the covariance functions is diagonal noise, the same problem emerges when combining additive covariance functions with Kronecker structure: the resulting covariance function components cannot be interpreted totally straightforwardly, as they are multiplied with the covariance functions associated with the other Kronecker components. Formally, a traditional random effects model is represented in the first half of the following inequality, and the model we implemented in the second:
$(\mathbf{K}_{1}\otimes\mathbf{K}_{2})+(\sigma^{2}\mathcal{I}_{1}\otimes\mathcal{I}_{2})\neq(\mathbf{K_{1}}+\sigma^{2}\mathcal{I}_{1})\otimes\mathbf{K}_{2}$ (9)
This formulation of a random effect is equivalent to adding some extra variance to the coefficients corresponding to non-patient specific covariates. We detail this in Appendix B, for the case where all kernels are linear and thus GP regression is equivalent to Bayesian linear regression.
We ran separate models with and without the "random effect" noise in the individual-level covariance function to explore its influence on the model fit and predictions. We call the models with and without this the mixed-effect GP ("GP.m") and the fixed-effect GP ("GP.f").
The covariance functions for each covariate were chosen conditional on the data. For continuous-valued covariates, a squared-exponential covariance function with a lengthscale hyperparameter enabling automatic relevance determination (ARD) was used. For binary covariates, a linear covariance function was used with a variance hyperparameter, as a more complex covariance function would be unnecessary for binary data. Nominal or ordinal covariates were one-hot encoded to corresponding binary variables and a linear covariance function used. The multi-output covariance matrix $\mathbf{K}_{out}$ is parametrised as a Cholesky-decomposed correlation matrix multiplied with a diagonal matrix representing the separate output variances.
2.6 Full Model Specification
In summary, the three-component Kronecker model used in this article, with $n_{g}$ Gaussian-distributed outputs $y_{g}$, $n_{b}$ Bernoulli-distributed outputs $y_{b}$, and repeated covariates partitioned into repeated measure grid components $x_{1}$, $x_{2}$ and $x_{3}$, becomes:
$\rho_{1},\rho_{2},\rho_{3}\sim\mathrm{InvGamma}(2,1)$ Lengthscales for each kernel Kronecker component
$\alpha\sim\mathrm{InvGamma}(2,1)$ Kernel variances for each output
$\alpha^{(n)}\sim\mathrm{InvGamma}(2,1)$ Noise variance for each Gaussian-distributed output
$\eta\sim\mathcal{N}(0,1)$ Standard Normal latent $\eta$ for computation purposes
$L\sim\mathrm{LkjCholesky}(3)$ Correlation matrix between outputs with LkjCholesky prior
$L^{(n)}\sim\mathrm{LkjCholesky}(3)$ Noise correlation matrix between Gaussian outputs with LkjCholesky prior
$\sigma^{2}_{re}\sim\mathrm{InvGamma}(2,1)$ Variance of optional random effects kernel component
$K_{1}=k(x_{1},x_{1},\rho_{1})$ First Kronecker kernel matrix component
$K_{2}=k(x_{2},x_{2},\rho_{2})$ Second Kronecker kernel matrix component
$K_{3}=k(x_{3},x_{3},\rho_{3})+\sigma^{2}_{re}\mathcal{I}$ Third Kronecker kernel matrix component, with optional random effects
$f=\mathrm{diag}(\alpha)L\otimes\mathrm{chol}(K_{3})\otimes\mathrm{chol}(K_{2})\otimes\mathrm{chol}(K_{1})\eta$ Latent variable $f$ constructed from kernels and $\eta$
$\Sigma^{(n)}=\mathrm{diag}(\alpha^{(n)})L^{(n)}(\mathrm{diag}(\alpha^{(n)})L^{(n)})^{T}$ Noise covariance between Gaussian-distributed outputs
$y_{g}\sim\mathcal{N}(f[1:n_{g},],\Sigma^{(n)})$ Distribution of $n_{g}$ Gaussian-distributed outputs $y_{g}$
$y_{b}\sim\mathrm{Bernoulli}(\Phi(f[n_{g}+1:n_{g}+n_{b},]))$ Distribution of $n_{b}$ Bernoulli-distributed outputs $y_{b}$
2.7 Stan Implementation
The methods described here were implemented in Stanstancore, a probabilistic programming language designed for Bayesian inference, in which model specification is performed explicitly within the language, and No U-Turn Sampler (NUTS) Hamiltonian Monte Carlohoffman2014no (HMC) is performed for the model parameters. The Gaussian Process was represented for inference with a standardised, uncorrelated Gaussian latent variable $\eta$, which was reshaped and transformed by the covariance Kronecker components to form the latent variable $f$ conditioned on the covariance. HMC was performed for the GP hyperparameters and noise terms, the latent function representation $\eta$, the latent function $f$ for discrete-valued tasks, and the missing values of the outputs. The Stan programs were called from within the R programming language via Rstanrstan. Similar computational speedups have been achieved in Stan previously flaxman2015fast, Ronneberg2021.
2.8 Missing Data
In the case of missing values of the output variable y, which often arise when exploiting Kronecker structure over an incomplete grid, the Bayesian framework offers a convenient solution: the missing y values are treated as parameters to be inferred, using the distribution implied by the latent variable $f$ and likelihood $p(y|f)$, conditional on the observed covariates. The scalability associated with the Kronecker decomposition is preserved, while interpretable posterior distributions for the unobserved data are provided.
The inference for continuous-valued missing outputs $y$ can be integrated in a straightforward way with most inference schemes. Given that Stan cannot perform inference over discrete variables, then the missing binary outputs pose a challenge. This was resolved by calculating the log-likelihood contributions according to the Bernoulli likelihood with an appropriate link function, with the missing $y$ varying continuously between zero and one. This allows for efficient gradient-informed inference procedures, and the interpretation of the imputed variables as probabilities.
2.9 Radiological Image Quality
The clinical theme of this article is radiological image quality, a field in which it is common to combine continuous-valued "objective" measurements of image quality with discrete-valued "subjective" expert evaluations of image quality. This article uses two real data sets, the first of which is observational time-series data, and the second of which comes from a RCT with thorough repeated measures. In each case the effect of primary interest is the effect on the image quality of volume of contrast medium per body weight.
The first real data set was previously published as a clinical study in tran2022indirect. The image quality was measured in venal blood vessels of 53 patients, with continuous-valued attenuation in Hounsfield Units as the objective endpoint, and binary evaluations of image quality from three different consultant radiologists as the subjective endpoints. Each patient was injected with a different quantity of contrast medium per body weight, and evaluations of the images were taken at six thirty-second interval time points after injection. Evaluations were also performed for each side of the body. Further covariates used were gender, age and tube voltage. Further clinical details can be found here. The combination of 53 patients, 4 endpoints, 7 time points and 2 body sides implies a covariance structure of size 2,968, but the use of the Kronecker product for each of these contributions means that the computations are readily tractable on a personal laptop.
The second real data set comes from a Randomised Control Trial assessing image quality in arterial blood vessels following Computed Tomography Angiography (CTA) kristiansen2023halved. 210 patients were randomised to receive full versus half doses of contrast volume per body weight, with measurements repeated at 7 different locations in the body, and 3 different spectral energy levels measured in kiloelectronVolts (keV). Examples of images recorded at different body levels and energy levels are presented in Figure 1. Attenuation level and image noise recorded in selected Regions Of Interest (ROIs), each measured in Hounsfield units were used as continuous-valued "objective" outcomes, and two consultant radiologists evaluated every image on a nominal scale for image quality. Because of the very skew distribution of the nominal data, these were simplified to a binary variable of "Excellent" vs every other level, results into two binary "subjective" endpoints. Examples of the ordinal scale of subjective image quality are shown in Figure 2. Sex, age and flow speed of injection were also collected as relevant covariates. The combination of 210 patients, 4 endpoints, 7 body locations and three energy levels resulted in a data set of size $N=17,640$ but again the Kronecker structure enabled exact inference with limited computational resources.
Figure 1: Lower extremity CTA in the control group and in the experimental group with energy levels at 50 and 40 keV. ROIs were placed in the abdominal aorta, common iliac artery, superficial femoral artery, and popliteal artery show attenuation and noise values. Reproduced with permission from kristiansen2023halved.
Figure 2: Examples of ordinal scale for subjective examination quality. Reproduced with permission from kristiansen2023halved.
3 Related Work
Gaussian Processes have been used for decades under various names such as kriging or Bayesian kernel regression, and are related to the widely used Support Vector Machines (SVMs). Their used has increased in the past couple of decades with the advent of greater computational resources, and much research has occurred concerning their scalability to larger data sets under contemporary computational constraints. These methods often include variational inference schemes via the "inducing point" frameworkhensman2013gaussian, or spectral methods to approximate the full covariance function via sampled Fourier featuresrahimi2007random. The speedups possible through Kronecker structure have been used for some time, but have previously been used on structured image datagilboa2014image, drug combinationsRonneberg2021, Ronneberg2023, spatio-temporal modellingGelfand2004, and multi-task regression bonilla2007multi.
Heterogeneous multiple outputs have also received research focus: while multiple Gausssian-distributed outputs allow for analytical marginalisation of the latent variable, the presence of discrete or other non-Gaussian-distributed outputs forces the use of non-conjugate inference methods such as variational methods, Expectation Propagation or sampling of the latent variable. Here we opt for the latter solution, aided by the development of the Stan programming language and the underlying continuity and smoothness of all latent parameters of interest.
4 Results
The results describing the behaviour of the trained models are presented here. We ran experiments using 10-fold cross-validation for two GP models with and without the random effect component in the individual-level covariance function ("GP.m" and "GP.f"), and two parametric models run in brms also with or without individual-level random effects ("brms.m" and "brms.f").
4.1 Comparison in brms
Two brms models were used as comparison methods: ("brms.m" and "brms.f") with or without random effects at the individual level, respectively. Both models had linear fixed effects based on the covariates for each data set, i.e. for the simulated data example, each output modelled with the following model formulae:
 form.yi <-  yi ~ (1|p|id) + x1 + x2 + x3 # "brms.m"
 form.yi <-   yi ~ x1 + x2 + x3 # "brms.f"
with $p$ being shared between the outputs. Either Gaussian or Binomial likelihoods were then added to the bf objects, and all outputs were learned jointly in a single call to brm. Code is available in the supplementary material.
For the first real data set, an interaction was included between time and contrast volume per unit body mass, and for the second real data set, an interaction was included between randomisation group and energy level.
4.2 Data Simulation
Two distinct latent parametric functions were used to simulate the data, providing a nonlinear relationship between the outputs and the covariates:
$f_{1}=\exp(.15x_{1})-.6x_{2}^{2}+\sin(3x_{3})\$ (10)
$f_{2}=-\exp(-.15x_{1})+\lvert 3x_{2}\rvert-\cos(3x_{3})$ (11)
Two further latent functions were defined for the binomial outputs: $f_{3}=-f_{1}$ and $f_{4}=f_{2}$. All four of the latent functions then had Gaussian noise of mean zero and standard deviation 0.1 added: the resulting noisy samples for the first two outputs became the observed continuous variables, while the final two noisy samples were pushed through a probit link function and rounded to generate the observed binomial variables. No random effects were included in the data simulation process.
The covariates were sampled from $\mathcal{U}_{[-5,5]}$. 20 unique grid points were sample for $x_{1}$, 7 for $x_{2}$, and 3 for $x_{3}$. Combined with the four outputs and the Kronecker structure, this resulted in a total number of 1,680 unique observations.
The statistical models that included random effects at the individual level ("GP.m" and "brms.m") treated the third covariate $x_{3}$ as representing a measurement at the individual-level data, for the sake of comparison, but this choice is not expected to have a large influence on the results, as there was no variation in the simulated data beyond the observed covariates and shared noise.
4.3 Losses and Testing
The models were evaluated using 10-fold cross-validation to predict the test outputs given the training data and test covariates. Loss functions were evaluated using the mean predictive $f_{pm}$ across posterior samples and the observed data, using an absolute/L1 loss for Gaussian outputs $y_{g}$ and for binomial outputs $y_{b}$, the logarithm of one minus the probability of the observed data ("the log probability of the wrong answer"), i.e. $l(y_{g},f_{pm})=|y_{g}-f_{pm}|$ and $l(y_{b},f_{pm})=\log(\Phi(-f_{pm}*(2y_{b}-1))$, where $\Phi$ is the probit link.
This resulted in distinct populations of losses per method, with one evaluated loss per output data point. These are plotted as histograms for each output in Figures 3, 4,5, with the binomial losses represented on the log scale as well as the probability scale. The differences between populations of losses for each output were tested formally using a paired Wilcoxon rank sum test, corrected for four-fold multiple testing degeneracy between methods. The p-values from the Wilcoxon tests were further supplemented by rank-biserial correlations as effect sizes. Nonparametric tests were chosen for model evaluation in order to avoid making distributional assumptions about the populations of losses, and to make the results robust to monotonic transformations in the losses.
For each of the four models on each of the data sets, we present posterior summaries of the interpretable model parameters from the first of the ten CV folds in tables 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14 and 15. Considering the size of the data sets and the randomisation process used, we consider posterior summaries of a single CV fold to be representative of the population as a whole.
4.4 Results for Simulated Data
We see the predictive loss results in Table 1 and Figure 3. We see that the brms model effectively fails and returns samples from the prior on binomial output 1 and Gaussian output 1, which are derived from the same latent parametric function. The GP models consequently registered significantly lower losses with large effect sizes. We see a small and possibly spurious effect of the GP models appearing to not predict around the prior at $p=0.5$. For the second continuous output, we again see the GP models outperforming the brms models, with the GP.f performing substantially better than GP.m. For the second binomial output, we see that the brms models get many of the labels correct with high confidence, but also many of the labels wrong with high confidence. Consequently, the results here are more mixed, with the only significant differences being brms.m outperforming both GP.m and brms.f. We would expect from the absolute value present in 11 that the smooth GP model would find the second and fourth endpoints more challenging to model accurately.
Parameter posterior summaries are presented for all four models trained on this data set in tables 4, 5, 6 and 7.
GAUSSIAN BINOMIAL
GP.m brms.f brms.m GP.f GP.m brms.f brms.m GP.f
1 GP.m - -1.000 -1.000 -0.102 - -0.778 -0.778 0.077
brms.f <.001 - 0.062 1.000 <.001 - -0.005 0.769
brms.m <.001 0.439 - 1.000 <.001 0.953 - 0.769
GP.f 0.405 <.001 $0$ 0.664 <.001 <.001 -
2 GP.m - -0.567 -0.567 0.588 - -0.083 0.251 -0.074
brms.f <.001 - 0.084 0.790 0.892 - 0.598 0.018
brms.m <.001 0.291 - 0.794 0.008 <.001 - -0.196
GP.f <.001 <.001 <.001 - 0.892 0.892 0.055 -
Table 1: Results for the simulated data example, comparing between populations of losses for each modelling method, with the first Gaussian output in the upper left, the first Binomial to the upper right, the second Gaussian to the lower left, and the second Binomial to the lower right. Within each square, the left-lower-triangular results are the p-values from each wilcoxon rank sum test, while the right-upper-triangular results are the rank-biserial correlations representing effect sizes.
Figure 3: Figures for the simulated data example, representing histograms of losses for each of the method for each of the outputs. The uppermost two plots represent the losses on the two Gaussian-distributed outputs. Binomial outputs are plotted in the lower two rows, with those on the left on the exponentiated probability scale, and those on the right the same results plotted on the original logarithmic scale.
4.5 Observational Time-Series Data
We see the predictive loss results for the first real data set in Table 2 and Figure 4. We see that all of the methods were confidently correct in their predictions of labels for the binomial outputs, suggesting that this is a relatively easy classification problem. That said, when considering the ranking of the predictions, we see that GP.m model outperforms the other methods on the binomial outputs. We see a small number of extremely confident predictions from the brms model, especially for output 3, which is a concern even if they are correct.
When considering the continuous output, we see more mixed results between the methods, with the only clearly significant result being that the GP.f model appears to perform uniformly worse than the other methods.
Parameter posterior summaries are presented for all four models trained on this data set in tables 8, 9, 10 and 11.
GP.m brms.f brms.m GP.f GP.m brms.f brms.m GP.f
G.1 GP.m - 0.057 0.064 -0.359 - -0.530 -0.435 -0.839 GP.m B.2
brms.f 0.356 - 0.134 -0.405 <.001 - 0.840 -0.061 brms.f
brms.m 0.356 0.015 - -0.408 <.001 <.001 - -0.442 brms.m
GP.f <.001 <.001 <.001 - <.001 0.200 <.001 - GP.f
B.1 GP.m - -0.941 -0.687 -0.750 - -0.482 -0.176 -0.879 GP.m B.3
brms.f <.001 - 0.955 0.390 $0$ 0.987 -0.010 brms.f
brms.m <.001 $0$0.098 0.000 <.001 - -0.686 brms.m
GP.f <.001 <.001 0.038 - <.001 0.835 <.001 - GP.f
Table 2: Results for the observational time series data set, comparing between populations of losses for each modelling method, with the Gaussian output in the upper-left, and the three Binomial outputs in the lower-left, upper-right, and lower-right. Within each square, the left-lower-triangular results are the p-values from each wilcoxon rank sum test, while the right-upper-triangular results are the rank-biserial correlations representing effect sizes.
Figure 4: Figures for the observational time series data set, representing histograms of losses for each of the method for each of the outputs. Binomial outputs are plotted in the first three rows, with those on the left on the exponentiated probability scale, and those on the right the same results plotted on the original logarithmic scale. The lowest-left plot represents the losses on the one continuous output
4.6 Randomised Control Trial Data
We see the predictive loss results for the second real data set in Table 3 and Figure 5. We again see that all four of the methods predict the correct label confidently most of the time, again suggesting that this is a relatively easy prediction problem. The brms.f method appears to return a small number of predictions from the prior for the second binomial output. We see from the hypothesis testing that the GP.m model outperforms the other three methods for the first binomial output (with a small effect size relative to GP.f), whereas the GP.f model outperforms the other methods for the second binomial output.
For the continuous outputs, the GP.m model outperforms the three other methods in prediction, while the GP.f model performs slightly better or equally well as the brms models.
Parameter posterior summaries are presented for all four models trained on this data set in tables 12, 13, 14 and 15.
GAUSSIAN BINOMIAL
GP.m brms.f brms.m GP.f GP.m brms.f brms.m GP.f
1 GP.m - -0.598 -0.599 -0.447 - -0.974 -0.776 -0.197
brms.f <.001 - -0.014 0.134 <.001 - 0.989 0.859
brms.m <.001 0.423 - 0.135 <.001 <.001 - 0.503
GP.f <.001 <.001 <.001 - <.001 <.001 <.001 -
2 GP.m - -0.358 -0.360 -0.356 - -0.940 -0.680 0.417
brms.f <.001 - -0.088 -0.008 <.001 - 0.881 0.817
brms.m <.001 <.001 - -0.005 <.001 <.001 - 0.643
GP.f <.001 1.000 1.000 - <.001 <.001 <.001 -
Table 3: Results for the RCT data set, comparing between populations of losses for each modelling method, with the Gaussian output in the upper-left, and the three Binomial outputs in the lower-left, upper-right, and lower-right. Within each square, the left-lower-triangular results are the p-values from each wilcoxon rank sum test, while the right-upper-triangular results are the rank-biserial correlations representing effect sizes.
Figure 5: Figures for the RCT data example, representing histograms of losses for each of the method for each of the outputs. Binomial outputs are plotted in the first two rows, with those on the left on the exponentiated probability scale, and those on the right the same results plotted on the original logarithmic scale. The lowest two plots represent the losses on the two Gaussian-distributed outputs.
5 Discussion
In this work, we have demonstrated that the repeated measurements used in applied medical research can be used to scale complex Bayesian nonparametric Gaussian Process models to practical clinical questions. The more complex models have showed to have increased predictive ability compared to the widely-used parametric models, indicating that this may be worthwhile, if achieving optimal model specification is a concern.
Model specification is particularly relevant in medical work when there are specific clinical questions or causal hypotheses to investigate: a poorly-specified model is prone to bias estimates of the parameters of interest and hence possibly give misleading results. The question remains as to how to represent clinical hypotheses within a GP framework: while in principle parametric regression models with finite linear combinations of features have corresponding covariance function representations, it may not be straightforward to interpret covariance function variances and lengthscales in a practical way. Preexisting familiarity with ANOVA and similar models that explain contributions to the variance rather than the functional form of the mean may be a helpful reference point.
For randomised data, in which one of the covariate dimensions has been generated by randomised interventions, it is common to estimate causal effects with parametric models, sometimes including other covariates to increase power. Such estimates will be potentially biased by the limited expressive capabilities of the parametric representation and corresponding model misspecification issues. If including other covariates, the covariance function representation would have the advantage of including more flexible function spaces that might more accurately reflect the true generative process, with interpretation of the parameters associated with non-randomised covariates being less important. We can therefore expect a more accurate estimation of the underlying causal effect when using a more flexible nonparametric model.
With non-randomised observational data, we may still be interested in achieving some insight into the causal mechanism underlying the data, with all of the appropriate caveats of potential confounding. In this case, the covariance function representation may still help by providing a more flexible function space that increases precision of the causal estimate of interest, and possibly reducing the bias of the estimated causal effect by more accurately modelling the influence of observed confounders.
The separability of the covariance functions necessary to exploit Kronecker structure is potentially an important limitation: it is somewhat analogous to being forced to include interaction effects in a parametric model. If the practitioner is interested in extracting a standalone main effect for the purposes of interpretation, or encoding separate mean functions as random effects at an individual level, then this may hinder interpretation. Further work exploring the interpretation of additively structured Kronecker sub-components would help to elucidate the implications of this constraint. As observed earlier, the stricter constraint of having the training data lie on a (mostly) complete grid is obeyed surprisingly frequently in clinical research, as the importance of performing repeated measurements to isolate different contributions to variation is well-understood.
The use of a Gaussian Process object also opens up the possibility of using Bayesian Optimizationfrazier2018tutorial type algorithms to measure new data points that are optimally informative, according to some acquisition function. The analytical posterior mean and variance of the latent GP object could be used to assess which unsampled areas of the covariate space would be best explored to reduce uncertainty in the effect of interest. Given the relatively strict grid structure necessary, performing this iteratively may be challenging, but the posterior predictive estimates provided by the fitted GPs could still be used to motivate future research study designs.
In conclusion, in this article, we have demonstrated the ability and utility of scaling Gaussian Process models to large real-world clinical data sets through the use of Kronecker-structure covariance matrices and repeated measurements in the data.
Acknowledgments
We thank Thien Trung Tran, Cathrine Helgestad Kristiansen and Peter Lauritzen for providing the radiological data used in this article.
Author contributions
OT helped with conceptualising the project, constructing the code, running the experiments, and writing the manuscript. LR helped with model design, providing the starting codebase, writing the manuscript, and deriving theoretical results.
Financial disclosure
None reported.
Conflict of interest
The authors declare no potential conflict of interests.
Appendix A Stan Outputs
mean sd l-95% CI u-95% CI n.eff Rhat
$\rho_{1}$ 1.404 0.173 1.087 1.771 321.774 1.006
$\rho_{2}$ 0.97 0.076 0.823 1.12 947.797 1.007
$\rho_{3}$ 1.193 0.112 0.959 1.408 202.334 1.009
$\alpha_{1}$ 0.982 0.143 0.752 1.31 623.841 1.003
$\alpha_{2}$ 1.497 0.273 1.048 2.113 304.171 1.007
$\alpha_{3}$ 1.005 0.279 0.575 1.652 1373.194 1.005
$\alpha_{4}$ 1.735 0.42 1.068 2.718 609.914 1.006
$\alpha^{(n)}_{1}$ 0.106 0.007 0.093 0.119 1916.695 1.002
$\alpha^{(n)}_{2}$ 0.132 0.009 0.115 0.149 650.408 1.001
$L^{(n)}_{21}$0.005 0.09 -0.185 0.169 2174.739 1.001
$L^{(n)}_{22}$ 0.996 0.006 0.98 1 2543.158 1
$L_{21}$0.135 0.143 -0.416 0.149 439.501 1.006
$L_{22}$ 0.98 0.025 0.909 1 476.384 1.007
$L_{31}$ 0.686 0.143 0.365 0.903 2331.114 1
$L_{32}$0.08 0.196 -0.432 0.319 2916.251 1
$L_{33}$ 0.667 0.137 0.393 0.912 1995.937 1.001
$L_{41}$0.034 0.163 -0.347 0.278 644.503 1.003
$L_{42}$ 0.867 0.085 0.646 0.976 2074.728 1
$L_{43}$0.022 0.184 -0.392 0.336 1443.446 1.002
$L_{44}$ 0.4 0.137 0.165 0.701 1730.865 1
Table 4: Posterior summaries of the interpretable parameters of the GP.f model from the first CV fold on the simulated data example from Section 4.4
mean sd l-95% CI u-95% CI n.eff Rhat
$\rho_{1}$ 1.833 0.171 1.507 2.188 630.434 1.004
$\rho_{2}$ 1.018 0.064 0.897 1.146 1156.583 1.004
$\rho_{3}$ 24.54 32.778 5.98 89.383 1883.495 1.003
$\alpha_{1}$ 0.746 0.095 0.579 0.954 889.352 1.003
$\alpha_{2}$ 2.076 0.261 1.613 2.624 577.406 1.007
$\alpha_{3}$ 0.863 0.264 0.48 1.487 1686.661 1.002
$\alpha_{4}$ 2.464 0.541 1.541 3.632 1534.739 1.003
$\alpha^{(n)}_{1}$ 0.104 0.007 0.092 0.118 2983.666 1.001
$\alpha^{(n)}_{2}$ 0.101 0.007 0.088 0.115 1881.657 1.001
$L^{(n)}_{21}$0.207 0.09 -0.383 -0.025 3693.566 1
$L^{(n)}_{22}$ 0.974 0.02 0.924 0.999 3455.251 0.999
$L_{21}$0.183 0.126 -0.416 0.076 364.228 1.01
$L_{22}$ 0.975 0.025 0.91 1 411.066 1.008
$L_{31}$ 0.605 0.164 0.226 0.865 2594.035 1.002
$L_{32}$0.046 0.221 -0.46 0.398 3936.939 1
$L_{33}$ 0.734 0.132 0.456 0.955 2001.942 1.003
$L_{41}$0.144 0.141 -0.405 0.147 599.615 1.006
$L_{42}$ 0.85 0.08 0.664 0.964 2309.363 1
$L_{43}$0.039 0.19 -0.424 0.324 1337.541 1.008
$L_{44}$ 0.421 0.127 0.184 0.675 2178.473 1
$\sigma^{2}_{re}$ 0.058 0.012 0.039 0.086 1102.067 1.004
Table 5: Posterior summaries of the interpretable parameters of the GP.m model from the first CV fold on the simulated data example from Section 4.4
Estimate sd l-95% CI u-95% CI Rhat Bulk.ESS Tail.ESS
y1.Intercept 0.002 0.068 -0.134 0.135 1.001 5464.966 3092.083
y2.Intercept -0.161 0.055 -0.268 -0.055 1.002 4786.035 2940.475
y3.Intercept 0.058 0.147 -0.234 0.345 1 5457.726 3301.1
y4.Intercept -0.265 0.173 -0.611 0.068 1 5162.581 2977.072
y1.x1 0.234 0.082 0.075 0.396 1.002 5932.14 3073.043
y1.x2 0.421 0.073 0.277 0.565 1 5867.37 2870.77
y1.x3 -0.03 0.078 -0.185 0.121 1 4867.733 2544.01
y2.x1 -0.49 0.067 -0.623 -0.362 1 4672.154 2914.568
y2.x2 0.446 0.059 0.335 0.559 1 4995.344 2801.483
y2.x3 0.359 0.062 0.24 0.477 1.001 4879.298 3101.912
y3.x1 0.043 0.182 -0.304 0.388 1.002 5430.894 3058.107
y3.x2 0.078 0.16 -0.23 0.386 1.001 6038.251 3178.676
y3.x3 -0.032 0.169 -0.36 0.301 1.001 5622.199 3016.555
y4.x1 -0.913 0.217 -1.349 -0.507 1 5067.848 3007.645
y4.x2 0.933 0.209 0.538 1.356 1.003 4827.501 3173.996
y4.x3 0.435 0.193 0.06 0.809 1 4120.251 3021.145
Table 6: Posterior summaries of the parameters of the brms.f model from the first CV fold on the simulated data example from Section 4.4
Estimate sd l-95% CI u-95% CI Rhat Bulk.ESS Tail.ESS
y1.Intercept 0.002 0.077 -0.151 0.152 1.001 6402.526 2842.214
y2.Intercept -0.15 0.286 -0.736 0.438 1.001 1398.978 1713.599
y3.Intercept 0.06 0.183 -0.297 0.418 1.001 4684.172 2849.119
y4.Intercept -0.288 0.546 -1.362 0.854 1.001 1799.947 2199.357
y1.x1 0.235 0.082 0.07 0.391 1 9351.66 2680.745
y1.x2 0.423 0.073 0.28 0.565 1.002 9552.251 3006.171
y1.x3 -0.028 0.09 -0.197 0.151 1.001 5891.755 3109.685
y2.x1 -0.49 0.029 -0.546 -0.434 1 8431.978 2622.935
y2.x2 0.445 0.025 0.397 0.492 1.001 8605.844 2679.603
y2.x3 0.35 0.324 -0.333 0.996 1.002 1828.779 2024.193
y3.x1 0.046 0.183 -0.315 0.399 1.001 7798.346 3147.908
y3.x2 0.078 0.155 -0.23 0.387 1.003 8106.717 2825.376
y3.x3 -0.03 0.219 -0.451 0.391 1 4265.744 2471.228
y4.x1 -1.154 0.257 -1.671 -0.668 1.001 6917.828 2769.571
y4.x2 1.182 0.257 0.712 1.701 1.003 8080.056 3051.224
y4.x3 0.577 0.614 -0.656 1.874 1.001 2099.679 2356.275
Table 7: Posterior summaries of the parameters of the brms.m model from the first CV fold on the simulated data example from Section 4.4
mean sd l-95% CI u-95% CI n.eff Rhat
Side 0.138 0.037 0.08 0.226 1514.041 1.002
Time 0.282 0.021 0.244 0.326 479.861 1.003
TubeVoltage 0.298 0.099 0.151 0.539 3460.95 1
Sex 1.057 0.218 0.689 1.545 2129.349 1.004
Age 0.32 0.092 0.217 0.602 75.444 1.026
KbyV 0.976 0.253 0.338 1.391 97.965 1.017
$\alpha_{1}$ 0.584 0.049 0.495 0.686 946.77 1.007
$\alpha_{2}$ 6.708 2.442 3.748 12.746 902.727 1.008
$\alpha_{3}$ 2.963 0.593 2.029 4.351 1385.233 1.006
$\alpha_{4}$ 5.118 3.198 2.613 11.731 233.638 1.017
$\alpha^{(n)}_{1}$ 0.447 0.017 0.415 0.481 1109.652 1.002
$L_{21}$ 0.767 0.072 0.607 0.884 705.094 1.002
$L_{22}$ 0.632 0.085 0.468 0.795 684.22 1.002
$L_{31}$ 0.813 0.06 0.677 0.91 1004.867 1.001
$L_{32}$ 0.512 0.091 0.338 0.691 974.675 1.001
$L_{33}$ 0.249 0.053 0.154 0.361 1884.588 0.999
$L_{41}$ 0.776 0.076 0.604 0.903 700.609 1.004
$L_{42}$ 0.486 0.114 0.26 0.695 693.147 1.004
$L_{43}$0.205 0.096 -0.386 -0.011 1136.005 1
$L_{44}$ 0.292 0.079 0.146 0.45 1506.603 1
Table 8: Posterior summaries of the interpretable parameters of the GP.f model from the first CV fold on the observational time series example from Section 4.5
mean sd l-95% CI u-95% CI n.eff Rhat
Side 0.811 0.171 0.527 1.205 1923.84 1.001
Time 0.257 0.017 0.225 0.292 507.519 1.005
TubeVoltage 0.249 0.097 0.117 0.48 3308.898 1
Sex 1.147 0.286 0.699 1.791 3691.531 1
Age 17.824 38.394 5.412 59.024 1914.216 1.002
KbyV 6.986 3.715 3.664 14.81 1589.906 1
$\alpha_{1}$ 0.424 0.052 0.335 0.536 1307.983 1.004
$\alpha_{2}$ 9.917 6.899 3.258 29.889 1739.195 1.002
$\alpha_{3}$ 6.743 4.765 2.485 20.149 973.293 1.001
$\alpha_{4}$ 6.568 5.372 1.985 22.38 1717.292 1.002
$\alpha^{(n)}_{1}$ 0.334 0.016 0.303 0.366 582.19 1.003
$L_{21}$ 0.242 0.089 0.064 0.415 2003.83 1.002
$L_{22}$ 0.966 0.023 0.91 0.998 2010.449 1.001
$L_{31}$ 0.354 0.085 0.18 0.511 1985.88 1
$L_{32}$ 0.848 0.045 0.748 0.926 1860.018 1.002
$L_{33}$ 0.377 0.065 0.256 0.514 1790.967 1.004
$L_{41}$ 0.211 0.099 0.016 0.402 2336.426 1
$L_{42}$ 0.785 0.065 0.639 0.891 2637.555 1.001
$L_{43}$0.33 0.121 -0.551 -0.078 1495.683 1.002
$L_{44}$ 0.438 0.107 0.226 0.64 1183.002 1.002
$\sigma^{2}_{re}$ 0.907 0.227 0.537 1.403 1096.364 1.001
Table 9: Posterior summaries of the interpretable parameters of the GP.m model from the first CV fold on the observational time series example from Section 4.5
Estimate sd l-95% CI u-95% CI Rhat Bulk.ESS Tail.ESS
HU.Intercept -1.39 0.074 -1.537 -1.247 1 2409.083 2984.331
Peter.Intercept -23.136 15.093 -62.433 -6.119 1.003 608.858 576.982
Sumit.Intercept -23.363 18.786 -77.043 -5.53 1.006 409.13 320.44
Thien.Intercept -29.083 19.913 -70.655 -7.379 1.003 691.947 498.643
HU.Sex -0.254 0.051 -0.357 -0.154 1.001 4559.867 2846.231
HU.Age -0.082 0.022 -0.125 -0.039 1.001 4361.788 2682.417
HU.KbyV 0.15 0.028 0.094 0.207 1.001 3720.045 3164.562
HU.TubeVoltage -0.007 0.061 -0.125 0.114 1 3883.938 3248.508
HU.Side -0.034 0.043 -0.12 0.05 1 4765.954 2693.787
HU.Times.1 0.772 0.08 0.615 0.926 1.001 2144.028 2807.039
HU.Times.2 1.699 0.082 1.537 1.866 1.001 2219.165 2908.287
HU.Times.3 2.093 0.083 1.932 2.262 1 2269.678 2760.094
HU.Times.4 2.161 0.08 2.003 2.318 1.001 2205.402 2842.623
HU.Times.5 2.16 0.081 1.998 2.32 1.001 2183.258 2429.128
HU.Times.6 2.159 0.081 2.003 2.317 1 2267.362 3046.82
Peter.Sex -1.596 0.392 -2.383 -0.842 1.001 3919.9 3341.905
Peter.Age -0.625 0.151 -0.931 -0.34 1.002 5417.444 2985.744
Peter.KbyV 0.296 0.19 -0.07 0.679 1 3325.893 2899.162
Peter.TubeVoltage 1.505 0.421 0.662 2.33 1.002 2808.877 2631.265
Peter.Side 0.101 0.283 -0.439 0.659 1 5410.548 2982.479
Peter.Times.1 19.842 15.096 2.695 58.589 1.003 609.021 557.101
Peter.Times.2 23.304 15.094 6.205 62.54 1.003 608.571 545.979
Peter.Times.3 25.541 15.108 8.451 64.9 1.003 607.598 567.206
Peter.Times.4 25.885 15.102 8.738 65.16 1.002 613.898 549.396
Peter.Times.5 26.028 15.118 8.906 65.348 1.003 611.03 581.704
Peter.Times.6 26.01 15.096 8.902 65.252 1.003 612.908 577.387
Sumit.Sex -1.364 0.325 -2.022 -0.746 1 4423.775 2895.507
Sumit.Age -0.551 0.127 -0.793 -0.303 1.001 5916.021 2742.276
Sumit.KbyV 0.5 0.154 0.205 0.808 1.001 3755.939 2738.673
Sumit.TubeVoltage 1.049 0.349 0.38 1.754 1.001 3319.896 2957.351
Sumit.Side -0.034 0.231 -0.495 0.415 1 5586.167 2965.208
Sumit.Times.1 20.573 18.783 2.707 74.057 1.006 408.385 317.588
Sumit.Times.2 23.32 18.79 5.438 76.846 1.006 408.887 320.442
Sumit.Times.3 25.238 18.792 7.402 79.115 1.006 409.265 319.583
Sumit.Times.4 25.295 18.788 7.376 78.9 1.005 409.093 321.446
Sumit.Times.5 25.529 18.783 7.659 78.986 1.005 409.948 320.44
Sumit.Times.6 24.87 18.781 7.055 78.473 1.006 409.119 321.843
Thien.Sex -1.093 0.516 -2.116 -0.103 1.001 3390.629 2920.511
Thien.Age -0.562 0.208 -0.988 -0.157 1 5525.843 2864.83
Thien.KbyV -0.035 0.276 -0.577 0.509 1 3796.754 3311.912
Thien.TubeVoltage 2.026 0.607 0.854 3.222 1 3172.331 2965.172
Thien.Side 0.195 0.405 -0.585 0.973 1.001 5236.692 2325.445
Thien.Times.1 24.506 19.934 2.887 66.063 1.003 691.319 479.517
Thien.Times.2 28.927 19.911 7.254 70.355 1.003 692.155 495.169
Thien.Times.3 32.396 19.948 10.587 74.065 1.003 689.624 489.518
Thien.Times.4 34.038 19.995 11.864 75.923 1.003 691.37 485.938
Thien.Times.5 34.031 19.989 11.909 76.187 1.003 695.89 504.097
Thien.Times.6 33.999 19.993 11.874 76.105 1.003 683.468 502.573
Table 10: Posterior summaries of the parameters of the brms.f model from the first CV fold on the observational time series example from Section 4.5
Estimate sd l-95% CI u-95% CI Rhat Bulk.ESS Tail.ESS
HU.Intercept -1.387 0.105 -1.591 -1.18 1.001 1618.397 2350.325
Peter.Intercept -33.07 17.474 -76.833 -11.798 1.001 610.753 403.092
Sumit.Intercept -29.19 28.557 -$1$8.163 1.033 96.551 22.218
Thien.Intercept -47.712 25.873 -119.681 -16.076 1.005 397.244 235.653
HU.Sex -0.249 0.104 -0.457 -0.045 1 1274.419 1887.114
HU.Age -0.083 0.045 -0.171 0.006 1.001 1298.837 1541.561
HU.KbyV 0.152 0.055 0.042 0.258 1.003 1179.928 2072.619
HU.TubeVoltage -0.003 0.12 -0.239 0.235 1.001 1153.473 1620.713
HU.Side -0.03 0.038 -0.104 0.044 1.007 8636.042 2727.371
HU.Times.1 0.772 0.069 0.636 0.906 1.001 3222.682 3355.595
HU.Times.2 1.701 0.071 1.563 1.837 1 3227.142 3142.511
HU.Times.3 2.092 0.072 1.95 2.233 1 3132.795 3264.571
HU.Times.4 2.16 0.07 2.019 2.301 1.001 3122.661 2327.567
HU.Times.5 2.16 0.07 2.024 2.303 1.001 3183.713 3003.711
HU.Times.6 2.15 0.072 2.007 2.29 1.001 3253.032 3026.83
Peter.Sex -4.558 1.801 -8.348 -1.288 1.001 1423.471 2296.226
Peter.Age -1.553 0.759 -3.166 -0.149 1 1302.35 2167.791
Peter.KbyV 0.033 0.909 -1.802 1.771 1.004 1102.953 1896.869
Peter.TubeVoltage 3.637 1.948 -0.117 7.662 1.001 1250.605 1622.67
Peter.Side 0.315 0.414 -0.483 1.116 1.001 8134.444 3079.231
Peter.Times.1 25.238 17.369 4.68 68.967 1.002 610.631 399.844
Peter.Times.2 34.946 17.552 13.457 79.709 1.001 603.05 408.878
Peter.Times.3 40.177 17.659 18.292 85.207 1.001 598.898 406.589
Peter.Times.4 40.714 17.667 18.804 85.767 1.001 600.787 407.753
Peter.Times.5 40.987 17.666 19.086 85.603 1.001 604.068 428.653
Peter.Times.6 40.957 17.673 19.018 85.762 1.001 612.928 424.599
Sumit.Sex -2.38 1.079 -4.695 -0.294 1.001 1356.179 1981.976
Sumit.Age -0.942 0.444 -1.834 -0.11 1.003 1391.182 2059.489
Sumit.KbyV 0.857 0.56 -0.212 1.956 1.003 1385.784 2115.462
Sumit.TubeVoltage 1.7 1.175 -0.576 4.029 1.002 1272.342 1796.238
Sumit.Side -0.081 0.301 -0.675 0.496 1.001 7380.028 2808.069
Sumit.Times.1 24.668 28.559 3.78 137.087 1.031 104.278 22.229
Sumit.Times.2 29.354 28.572 8.362 141.328 1.031 103.682 22.201
Sumit.Times.3 32.853 28.572 11.76 144.49 1.032 103.688 22.119
Sumit.Times.4 32.896 28.569 11.864 144.481 1.031 103.616 22.181
Sumit.Times.5 33.322 28.568 12.247 144.342 1.032 103.926 22.156
Sumit.Times.6 32.2 28.571 11.09 144.183 1.032 103.205 22.143
Thien.Sex -2.909 2.095 -7.371 0.794 1.001 1832.914 2101.764
Thien.Age -1.337 0.879 -3.253 0.247 1.001 1666.775 2237.486
Thien.KbyV -0.32 1.08 -2.498 1.775 1.004 1420.407 2315.322
Thien.TubeVoltage 5.034 2.561 0.484 10.854 1 1559.872 1907.257
Thien.Side 0.637 0.608 -0.522 1.839 1.003 6508.089 2826.776
Thien.Times.1 36.157 25.34 5.81 108.251 1.004 402.642 236.309
Thien.Times.2 48.473 25.996 16.357 121.175 1.004 391.273 239.821
Thien.Times.3 55.615 26.395 22.403 128.826 1.004 389.505 234.506
Thien.Times.4 57.812 26.517 24.452 130.267 1.005 390.179 235.505
Thien.Times.5 57.797 26.51 24.464 130.758 1.004 393.287 238.189
Thien.Times.6 57.725 26.516 24.307 131.551 1.004 391.34 230.294
Table 11: Posterior summaries of the parameters of the brms.m model from the first CV fold on the observational time series example from Section 4.5
mean sd l-95% CI u-95% CI n.eff Rhat
Location.Iliac 1.286 0.074 1.146 1.439 1513.542 1.003
Location.Femoral 9.77 1.739 7.113 14.038 445.735 1.005
Location.Popliteal 14.298 4.277 8.887 23.962 561.439 1.003
keV.40 0.237 0.026 0.187 0.29 759.765 1.002
keV.50 0.113 0.014 0.088 0.142 744.719 1.002
Randomisation 0.292 0.038 0.228 0.37 633.115 1.009
Sex 1.052 0.078 0.905 1.212 1591.639 1.007
Age 0.237 0.015 0.202 0.265 44.326 1.092
Flow 0.348 0.03 0.305 0.439 38.1 1.116
$\alpha_{1}$ 1.012 0.054 0.914 1.128 480.564 1.003
$\alpha_{2}$ 0.619 0.035 0.555 0.691 758.82 1.001
$\alpha_{3}$ 4.936 0.776 3.694 6.722 1065.099 1.002
$\alpha_{4}$ 15.009 2.419 11.154 20.61 1338.5 1
$\alpha^{(n)}_{1}$ 0.294 0.004 0.286 0.303 1306.745 1.003
$\alpha^{(n)}_{2}$ 0.715 0.009 0.697 0.733 1576.561 1.001
$L^{(n)}_{21}$0.144 0.019 -0.181 -0.107 2935.456 1.001
$L^{(n)}_{22}$ 0.989 0.003 0.984 0.994 2885.614 1.001
$L_{21}$ 0.732 0.033 0.661 0.79 694.7 1.003
$L_{22}$ 0.68 0.035 0.613 0.75 700.848 1.003
$L_{31}$ 0.602 0.044 0.512 0.685 1118.582 1.003
$L_{32}$0.032 0.073 -0.176 0.114 905.705 1.007
$L_{33}$ 0.793 0.034 0.722 0.857 1132.266 1.002
$L_{41}$ 0.71 0.031 0.647 0.767 985.695 1.003
$L_{42}$0.015 0.061 -0.135 0.104 492.138 1.006
$L_{43}$ 0.321 0.057 0.205 0.429 511.796 1.012
$L_{44}$ 0.619 0.038 0.545 0.693 628.993 1.005
Table 12: Posterior summaries of the interpretable parameters of the GP.f model from the first CV fold on the RCT example from Section 4.6
mean sd l-95% CI u-95% CI n.eff Rhat
Location.Iliac 1.071 0.095 0.9 1.275 496.725 1.008
Location.Femoral 1.69 0.112 1.482 1.931 718.188 1.002
Location.Popliteal 1.776 0.125 1.552 2.039 640.221 1.005
keV.40 0.102 0.012 0.079 0.127 1185.38 1.002
keV.50 0.054 0.007 0.041 0.068 1123.026 1
Randomisation 0.12 0.026 0.078 0.18 1160.001 1
Sex 1.31 0.166 1.022 1.668 2934.555 1
Age 35.694 12.951 19.5 69.44 812.477 1.004
Flow 11.488 2.314 7.528 16.652 471.008 1.001
$\alpha_{1}$ 1 0.077 0.861 1.16 691.075 1.006
$\alpha_{2}$ 1.317 0.113 1.11 1.548 632.407 1.006
$\alpha_{3}$ 8.796 1.699 6.114 12.616 909.123 1.005
$\alpha_{4}$ 14.696 3.03 10.203 21.863 877.625 1.005
$\alpha^{(n)}_{1}$ 0.243 0.003 0.237 0.249 2819.81 1
$\alpha^{(n)}_{2}$ 0.567 0.008 0.552 0.582 1474.628 1
$L^{(n)}_{21}$0.12 0.018 -0.158 -0.085 2945.145 1
$L^{(n)}_{22}$ 0.993 0.002 0.988 0.996 2900.402 1
$L_{21}$ 0.295 0.041 0.217 0.374 770.558 1.004
$L_{22}$ 0.955 0.013 0.927 0.976 768.216 1.004
$L_{31}$ 0.46 0.046 0.366 0.546 1245.77 1.002
$L_{32}$0.282 0.071 -0.416 -0.137 1191.492 1.003
$L_{33}$ 0.837 0.033 0.771 0.898 1171.472 1.003
$L_{41}$ 0.377 0.042 0.292 0.458 978.818 1.001
$L_{42}$0.053 0.061 -0.169 0.067 932.93 1.005
$L_{43}$ 0.106 0.079 -0.046 0.262 277.257 1.011
$L_{44}$ 0.912 0.02 0.871 0.948 907.221 1.003
$\sigma^{2}_{re}$ 0.121 0.018 0.09 0.159 769.687 1.003
Table 13: Posterior summaries of the interpretable parameters of the GP.m model from the first CV fold on the RCT example from Section 4.6
Estimate sd l-95% CI u-95% CI Rhat Bulk.ESS Tail.ESS
HU.Intercept 0.275 0.04 0.196 0.357 1 10453.102 2692.506
SD.Intercept -0.368 0.048 -0.462 -0.273 1.002 8602.66 3274.616
Sumit.Intercept 2.844 0.224 2.408 3.296 1.001 7851.73 3100.73
DanLevi.Intercept 2.046 0.158 1.735 2.343 1 7434.289 2993.31
HU.Sex -0.043 0.029 -0.1 0.014 0.999 7290.619 3315.779
HU.Age 0.076 0.013 0.051 0.101 1.001 8695.496 2708.414
HU.Flowmls 0.149 0.015 0.119 0.178 1 7565.725 3049.803
HU.Side 0 0.027 -0.05 0.054 1.001 8424.888 3120.189
HU.Location -0.085 0.013 -0.11 -0.059 1.001 9835.406 2940.605
HU.keV -0.625 0.013 -0.65 -0.6 1.005 9179.524 2829.488
SD.Sex -0.272 0.034 -0.339 -0.204 1.001 7248.65 3262.452
SD.Age -0.021 0.015 -0.051 0.009 1 8169.73 3188.157
SD.Flowmls 0.226 0.018 0.192 0.261 1.003 7449.155 3557.409
SD.Side 0.091 0.032 0.028 0.156 1.001 8420.121 3258.472
SD.Location 0.183 0.015 0.153 0.213 1.001 9028.951 2844.361
SD.keV -0.271 0.015 -0.301 -0.241 1.001 10044.46 2975.49
Sumit.Sex 0.053 0.16 -0.268 0.357 1 7327.927 3328.309
Sumit.Age -0.088 0.073 -0.234 0.049 1 10120.038 3351.897
Sumit.Flowmls 0.067 0.081 -0.093 0.22 1 6514.66 3419.375
Sumit.Side 0.139 0.148 -0.152 0.432 1.002 8806.937 2611.821
Sumit.Location 0.006 0.068 -0.127 0.141 1.002 8812.139 2832.606
Sumit.keV -0.847 0.069 -0.988 -0.717 1.001 6880.567 3395.213
DanLevi.Sex 0.204 0.115 -0.024 0.43 1.001 7937.028 2831.09
DanLevi.Age 0.02 0.05 -0.078 0.115 1.002 9383.669 2899.218
DanLevi.Flowmls 0.194 0.057 0.083 0.308 1 7261.622 3446.642
DanLevi.Side 0.042 0.101 -0.153 0.244 1 8716.294 3142.675
DanLevi.Location -0.141 0.047 -0.231 -0.05 1.002 7775.343 2986.236
DanLevi.keV -1.296 0.05 -1.396 -1.197 1 6444.788 3221.858
Table 14: Posterior summaries of the parameters of the brms.f model from the first CV fold on the RCT example from Section 4.6
Estimate sd l-95% CI u-95% CI Rhat Bulk.ESS Tail.ESS
HU.Intercept 0.238 0.088 0.062 0.401 1.026 187.541 469.566
SD.Intercept -0.386 0.075 -0.535 -0.24 1.005 418.06 1129.855
Sumit.Intercept 5.513 0.645 4.332 6.825 1.006 724.949 1647.996
DanLevi.Intercept 5.439 0.802 3.937 7.059 1.009 402.24 958.231
HU.Sex -0.024 0.121 -0.247 0.218 1.028 173.245 268.353
HU.Age 0.078 0.049 -0.017 0.174 1.019 255.92 531.007
HU.Flowmls 0.152 0.059 0.039 0.267 1.016 210.331 380.7
HU.Side 0.004 0.014 -0.024 0.032 1 5957.802 2980.97
HU.Location -0.075 0.007 -0.088 -0.062 1.002 6749.447 2902.081
HU.keV -0.626 0.007 -0.639 -0.612 1.001 8216.338 2970.541
SD.Sex -0.268 0.092 -0.438 -0.084 1.009 305.963 568.382
SD.Age -0.024 0.037 -0.097 0.048 1.011 408.812 947.47
SD.Flowmls 0.222 0.044 0.135 0.308 1.006 347.06 930.257
SD.Side 0.089 0.029 0.032 0.145 1.002 5594.598 2939.804
SD.Location 0.19 0.013 0.164 0.216 1.001 6120.002 3129.237
SD.keV -0.272 0.013 -0.297 -0.246 1.001 6320.706 2932.901
Sumit.Sex 0.267 0.698 -1.066 1.655 1.008 347.219 766.086
Sumit.Age -0.128 0.313 -0.771 0.46 1.012 502.922 1384.839
Sumit.Flowmls -0.034 0.356 -0.761 0.64 1.004 405.3 786.352
Sumit.Side 0.25 0.193 -0.132 0.636 1.002 4853.972 2810.804
Sumit.Location 0.006 0.085 -0.158 0.172 1.001 5430.007 3161.704
Sumit.keV -1.257 0.095 -1.451 -1.077 1.001 5131.408 3045.637
DanLevi.Sex 0.803 0.974 -1.141 2.675 1.008 305.69 595.933
DanLevi.Age -0.169 0.426 -1.004 0.67 1.007 462.015 1039.884
DanLevi.Flowmls 0.285 0.496 -0.7 1.238 1.002 410.709 1011.654
DanLevi.Side 0.14 0.19 -0.231 0.503 1 5126.812 3077.509
DanLevi.Location -0.39 0.091 -0.565 -0.21 1 4979.411 3335.896
DanLevi.keV -3.116 0.136 -3.389 -2.86 1 3505.023 2501.601
Table 15: Posterior summaries of the parameters of the brms.m model from the first CV fold on the RCT example from Section 4.6
Appendix B Random effects in the Intrinsic Coregionalization Model
In this appendix, we demonstrate the effect of adding a random effect to an Intrinsic Coregionalization Model (ICM), in the simplest case where the covariance functions over the inputs and the outputs are both linear. Recall that the ICM for a multi-output GP over $m$ outputs can be written as
$\begin{bmatrix}f_{1}\\ \vdots\\ f_{m}\end{bmatrix}\sim\mathcal{GP}(0,B\ \kappa_{x}(x,x^{\prime})),$ (12)
where $B$ is the $m\times m$ coregionalisation matrix giving the covariance over the outputs, and $\kappa_{x}(x,x^{\prime})$ the covariance function over the inputs. In the case of a fully observed dataset, we can write this using the Kronecker product $\mathbf{f}(\mathbf{X})\sim\mathcal{N}(0,B\otimes K_{x})$. By exploiting the connections between Gaussian Processes and Bayesian linear regression, in the case of a linear kernel over the inputs, $\kappa_{x}(\mathbf{x},\mathbf{x}^{\prime})=1+\mathbf{x}^{T}\mathbf{x}^{\prime}$, the ICM is equivalent to the following linear model:
$\begin{bmatrix}\mathbf{f}_{1}(\mathbf{X})\\ \vdots\\ \mathbf{f}_{m}(\mathbf{X})\end{bmatrix}=\begin{bmatrix}\mathbf{X}&0&\cdots&0\\ 0&\mathbf{X}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\mathbf{X}\end{bmatrix}\begin{bmatrix}\boldsymbol{\beta}_{1}\\ \boldsymbol{\beta}_{2}\\ \vdots\\ \boldsymbol{\beta_{m}}\end{bmatrix},\ \ \begin{bmatrix}\boldsymbol{\beta}_{1}\\ \boldsymbol{\beta}_{2}\\ \vdots\\ \boldsymbol{\beta_{m}}\end{bmatrix}\sim\mathcal{N}(0,B\otimes\mathcal{I}_{(p+1)}),$ (13)
where $\mathbf{X}\in\mathbb{R}^{m\times(p+1)}$ denotes the design matrix of a linear regression including the intercept, $\boldsymbol{\beta_{j}}\in\mathbb{R^{(p+1)}}$ the corresponding output-specific coefficient vector, and $\mathcal{I}_{p}$ the $(p+1)$-dimensional identity matrix. Hence, this is equivalent to fitting a linear model to each output, with output-specific coefficients $\boldsymbol{\beta}_{j}$ for $j=1,\ldots,m$. Due to the joint distribution over the coefficient vectors, these linear models are correlated in the prior, allowing the models to influence each other, borrowing strength across outputs.
When working with multiple covariance functions, that are multiplied together, we need to consider the effect on the induced feature space. Consider a set of covariates that can be blocked into two sets $\mathbf{X}=[\mathbf{X}^{(1)},\mathbf{X}^{(2)}]$, where $\mathbf{X}^{(2)}$ correspond to patient-specific covariates, while $\mathbf{X}^{(1)}$ are other covariates in the model. In this paper, we consider covariance structures that decompose across these groups of covariates, e.g. $\kappa(\mathbf{X},\mathbf{X}^{\prime})=\kappa_{1}(\mathbf{X}^{(1)},\mathbf{X}^{(1)})\kappa_{2}(\mathbf{X}^{(2)},\mathbf{X}^{(2)})$. Generally, multiplying together kernels has the effect of modelling interactions between the covariates, e.g. if both $\kappa_{1}$ and $\kappa_{2}$ are linear kernels, the induced features space is $\phi_{1\times 2}=\{1,x^{(1)}_{1},\ldots,x^{(1)}_{p_{1}},x^{(2)}_{1},\ldots,x^{(2)}_{p_{2}},x^{(1)}_{1}x^{(2)}_{1},\ldots,x^{(1)}_{p_{1}}x^{(2)}_{p_{2}}\}$, and GP regression is equivalent to performing a linear regression using the extended basis $\phi_{1\times 2}$. Similarly in the multi-output setting GP regression becomes equivalent to the model in equation (13), but with the $m\times((p_{1}+1)(p_{2}+1))$ matrix $\Phi_{1\times 2}$ taking the place of $X$, and modulo some dimensionality changes on the identity matrix in the prior over the coefficients.
Further complicating the model by adding a random effect to $\kappa_{2}$ as in Section 2.5:
$\begin{bmatrix}f_{1}\\ \vdots\\ f_{m}\end{bmatrix}\sim\mathcal{GP}\left(0,B\ \kappa_{1}\left(\mathbf{X}^{(1)},\mathbf{X}^{(1)}\right)\left(\kappa_{2}\left(\mathbf{X}^{(2)},\mathbf{X}^{(2)}\right)+\gamma^{2}\mathcal{I}_{m}\right)\right),$ (14)
has the effect of adding extra variance to the coefficients corresponding to the non-patient specific covariates. The model becomes
$\begin{bmatrix}\mathbf{f}_{1}(\mathbf{X})\\ \vdots\\ \mathbf{f}_{m}(\mathbf{X})\end{bmatrix}=\begin{bmatrix}\Phi_{1\times 2}&0&\cdots&0\\ 0&\Phi_{1\times 2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\Phi_{1\times 2}\end{bmatrix}\begin{bmatrix}\boldsymbol{\beta}_{1}\\ \boldsymbol{\beta}_{2}\\ \vdots\\ \boldsymbol{\beta_{m}}\end{bmatrix},\ \ \ \text{where}\ \ \ \begin{bmatrix}\boldsymbol{\beta}_{1}\\ \boldsymbol{\beta}_{2}\\ \vdots\\ \boldsymbol{\beta_{m}}\end{bmatrix}\sim\mathcal{N}\left(0,B\otimes\tilde{\mathcal{I}}_{(p_{1}+1)(p_{2}+1)}\right),$ (15)
where ~$\tilde{\mathcal{I}}_{(p_{1}+1)(p_{2}+1)}$ is a block diagonal matrix with blocks $\{(1+\gamma^{2})\mathcal{I}_{(p_{1}+1)},\mathcal{I}_{p_{2}},\mathcal{I}_{p_{1}p_{2}}\}$.